What is the y-coordinate of the vertex of the parabola y = -1.5(x - 2)2 + 7
Answer by lwsshak3(11628)

What is the y-coordinate of the vertex of the parabola y = -1.5(x - 2)2 + 7
..
y = -1.5(x - 2)2 + 7
This equation if a parabola of the standard form, y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
For given equation, y=-1.5(x-2)^2+7 is a parabola that opens downward with coordinates of the vertex at (2,7)
ans:
y-coordinate of vertex=7